# Burnside type results for fusion categories

### v10

@@ 3. Galois symmetries for hypergroups
Let $(H,\mathcal{B})$ be an abelian rational hypergroup. For any standard element $x_{i}\in\mathcal{B}$, one can write $x_{i}=\sum_{j=1}^{m}{\alpha_{{}_{ij}}}F_{j}$, for some scalars ${\alpha_{{}_{ij}}}\in\mathbb{C}$. Note that ${\alpha_{{}_{ij}}}=\mu_{j}(x_{i})$. Let $\mathbb{K}=\mathbb{Q}({\alpha_{{}_{ij}}})$ be the field obtained by adjoining all ${\alpha_{{}_{ij}}}\in\mathbb{C}$ to $\mathbb{Q}$.
@@ 3.1. Permutation of characters $\mu_{j}$
Let ${(H,\mathcal{B})}$ be any abelian rational hypergroup (i.e. $N^{k}_{ij}\in\mathbb{Q}$). For any character $\mu_{j}:H\rightarrow\mathbb{C}$ and $\sigma\in\mathrm{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$, define $\sigma.\mu_{j}\in H^{*}$ as the linear function on $H$ on which the standard basis $\mathcal{B}$ is given by $[\sigma.\mu_{j}](x_{i})=\sigma(\mu_{j}(x_{i}))=\sigma({\alpha_{{}_{ij}}}).$
@@
For any abelian rational hypergroup ${(H,\mathcal{B})}$, it follows that $\sigma.\mu_{j}:H\rightarrow\mathbb{C}$ is an algebra map. Thus $\mathrm{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$ acts on the set $^{\widehat{\mathcal{B}}}$.
@@
Indeed, suppose that $x_{i_{1}}x_{i_{2}}=\sum_{k=1}^{m}N^{k}_{{i_{1}},{i_{2}}}x_{k}$. Since $N^{k}_{{i_{1}},{i_{2}}}\in\mathbb{Q}$,
$[\sigma.\mu_{j}]({x_{i_{1}}}{x_{i_{2}}})=[\sigma.\mu_{j}](\sum_{k=1}^{m}N^{k}_{{i_{1}},{i_{2}}}x_{k})=\sum_{k=1}^{m}N^{k}_{{i_{1}},{i_{2}}}[\sigma.\mu_{j}](x_{k})=\sum_{k=1}^{m}N^{k}_{{i_{1}},{i_{2}}}\sigma(\mu_{j}(x_{k})).$
On the other hand, note that
@@
$=$ $\sigma(\mu_{j}({x_{i_{1}}}{x_{i_{2}}}))=\sigma(\mu_{j}(\sum_{k=1}^{m}N^{k}_{{i_{1}},{i_{2}}}x_{k}))$
$=$ $\sum_{k=1}^{m}N^{k}_{{i_{1}},{i_{2}}}\sigma(\mu_{j}(x_{k})).$
Thus $[\sigma.\mu_{j}]({x_{i_{1}}}{x_{i_{2}}})=[\sigma.\mu_{j}]({x_{i_{1}}})[\sigma.\mu_{j}]({x_{i_{2}}})$, so that $\sigma.\mu_{j}$ is an algebra map. It is easy to see that, in this case, $\sigma.(\sigma^{\prime}.\mu_{j})=(\sigma\sigma^{\prime}).\mu_{j}.$ It follows that $\mathrm{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$ acts on the set $^{\widehat{\mathcal{B}}}$ of all algebra homomorphisms $\mu_{j}:H\rightarrow\mathbb{C}$. ∎
We denote by $\mathcal{J}:=\{1,\dots,m\}$ the set of all indices for the characters $\mu_{j}:H\rightarrow\mathbb{C}$. It follows that for any $j\in\mathcal{J}$, there is an index $\tau(j)=\tau_{\sigma}(j)\in\mathcal{J}$ such that $\sigma.\mu_{j}=\mu_{\tau_{\sigma}(j)}$, and therefore
(3.1) $\sigma({\alpha_{{}_{ij}}})=\alpha_{{}_{i\tau_{\sigma}(j)}}\;\;\text{for all}\;i,j.$
It is easy to see that, in this case, $\tau_{\sigma}$ is a permutation of $\mathcal{J}$.
@@
Note that Equation (3.1) shows that $\sigma(\mathbb{K})\subseteq\mathbb{K}$, for all $\sigma\in\mathrm{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$. Therefore, this is a normal extension and therefore a Galois extension. ∎
@@
Note that the map $\rho:\mathrm{Gal}(\mathbb{K}/\mathbb{Q})\rightarrow\mathbb{S}_{m},\sigma\mapsto\tau_{\sigma}$ is injective. Indeed, if $\tau_{\sigma}=\tau_{\sigma^{\prime}}$, then $\sigma({\alpha_{{}_{ij}}})=\sigma^{\prime}({\alpha_{{}_{ij}}})$, i.e. $\sigma^{-1}\circ\sigma^{\prime}({\alpha_{{}_{ij}}})={\alpha_{{}_{ij}}}$, for all $i,j$. Thus $\sigma=\sigma^{\prime}$ on $\mathbb{K}$.
@@
Let $(H,\mathcal{B})$ be an abelian rational hypergroup. For any $\sigma\in\mathrm{Gal}(\mathbb{K}/\mathbb{Q})$, with the above notations,
(3.2) $\sigma(n_{k})={n_{\tau(k)}}.$
@@
Applying $\sigma\in\mathrm{Gal}(\mathbb{K}/\mathbb{Q})$ to Equation (2.5), one obtains the following:
$\sum_{i=1}^{m}{h_{i}}\sigma(\alpha_{{}_{il}})\sigma(\alpha_{{}_{i^{*}k}})=\delta_{l,k}\sigma(n_{k}).$
In turn, this equation can be written as
$\sum_{i=1}^{m}{h_{i}}\alpha_{{}_{i\tau(l)}}\alpha_{{}_{i^{*}\tau(k)}}=\delta_{l,k}\sigma(n_{k}).$
@@
$\sum_{i=1}^{m}{h_{i}}\alpha_{{}_{i\tau(l)}}\alpha_{{}_{i^{*}\tau(k)}}=\delta_{\tau(l),\tau(k)}{n_{\tau(k)}}.$
Therefore, for $l=k$, it follows that $\sigma(n_{k})={n_{\tau(k)}}.$
We continue with the following definition.
@@
A hypergroup ${(H,\mathcal{B})}$ is $h$-integral if $h_{i}:=1/N^{1}_{i,i^{*}}\in\mathbb{Z}$, for all $1\leq i\leq m$.
@@
In view of Equation (2.11), the Grothendieck ring of any weakly-integral fusion category $\mathcal{C}$ has a $h$-integral dual ${\widehat{{K(\mathcal{C})}}}$, see [Ost15, Theorem 2.13]. See also Remark 9.1.
@@
Suppose that ${(H,\mathcal{B},\mu_{1})}$ is an abelian normalizable hypergroup with a $h$-integral dual. Then ${(H,\mathcal{B},\mu_{1})}$ is weakly-integral. Moreover, if ${(H,\mathcal{B},\mu_{1})}$ is rational, then $\widehat{h}_{k}=\widehat{h}_{\tau(k)}$, for all $1\leq k\leq m$.
@@
By definition,
$n{(H,\mathcal{B},\mu_{1})}=n{\widehat{{(H,\mathcal{B},\mu_{1})}}}=\sum_{j=1}^{m}{\widehat{h}}_{j}\in\mathbb{Z}.$
If ${(H,\mathcal{B},\mu_{1})}$ is rational, then Equation (3.2) can be written as $\sigma\bigg{(}\frac{n{(H,\mathcal{B},\mu_{1})}}{\widehat{h_{k}}}\bigg{)}=\frac{n{(H,\mathcal{B},\mu_{1})}}{\widehat{h_{\tau(k)}}}.$ In particular, $\widehat{h}_{k}=\widehat{h}_{\tau(k)}.$, since ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is $h$-integral. ∎
@@ 4. Burnside’s vanishing property for hypergroups
Let $\mathcal{B}_{0}$ be the set of all vanishing elements of $\mathcal{B}$, and $\mathcal{B}_{1}$ be the set of all non-vanishing elements of $\mathcal{B}$. Thus $\mathcal{B}=\mathcal{B}_{1}\sqcup\mathcal{B}_{0}$. Note that if ${(H,\mathcal{B})}$ is a RN-hypergroup then ${G{(H,\mathcal{B})}}\subseteq\mathcal{B}_{1}$. Note that, in the case of an abelian hypergroup ${(H,\mathcal{B})}$, since $x_{i}=\sum_{j=1}^{m}\mu_{j}(x_{i})F_{j}$, it follows that $x_{i}$ is non-vanishing if and only if it is invertible. Thus, in this case, $\mathcal{B}_{1}$ coincides with the set of all standard elements that are invertible in $H$.
Denote ${\widehat{P}}:=\prod_{j=1}^{m}\mu_{j}$. Note that in any abelian normalizable ${(H,\mathcal{B},\mu_{1})}$ one has:
(4.1) $\prod_{j=1}^{m}\mu_{j}=\sum_{i=1}^{m}{\widehat{P}}({\frac{x_{i}}{d_{i}}}){\widehat{E}}_{i}.$
@@
Let ${(H,\mathcal{B},\mu_{1})}$ be any abelian normalizable hypergroup. Then the following holds:
(4.2) $\prod_{j=1}^{m}\mu_{j}=\sum_{x_{i}\in\mathcal{B}_{1}}\det(L_{{\frac{x_{i}}{d_{i}}}}){\widehat{E}}_{i}$
@@
Recall that $<{\widehat{E}}_{i},\frac{x_{j}}{d_{j}}>=\delta_{i,j}$ for all $i,j\in\mathcal{I}$. In order to prove the above equality, it is enough to evaluate both sides at each ${\frac{x_{i}}{d_{i}}}$, and prove that one obtains the same value.
If $x_{i}$ is a vanishing element, then the above left and right hand side both vanish when evaluated at ${\frac{x_{i}}{d_{i}}}$.
If $x_{i}$ is not a vanishing element then
$\big{[}\prod_{j=1}^{m}\mu_{j}\big{]}({\frac{x_{i}}{d_{i}}})=\prod_{j=1}^{m}\mu_{j}({\frac{x_{i}}{d_{i}}})=\det(L_{{\frac{x_{i}}{d_{i}}}})$
and the proof is compete. ∎
Dually, applying the above result to ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$, one obtains a decomposition of the basis ^=^${\widehat{\mathcal{B}}}={\widehat{\mathcal{B}}}_{0}\sqcup{\widehat{\mathcal{B}}}_{1}$. The above proposition implies the following:
@@
Let ${(H,\mathcal{B},\mu_{1})}$ be any abelian normalizable hypergroup. Then the following holds:
(4.3) $\prod_{i=1}^{m}{\frac{x_{i}}{d_{i}}}=\sum_{\mu_{j}\in{\widehat{\mathcal{B}}}_{1}}\det(L_{\mu_{j}})F_{j},$
@@
An abelian hypergroup ${(H,\mathcal{B})}$ is Burnside if and only if one has ${G{(H,\mathcal{B})}}=\mathcal{B}_{1}$.
@@
An abelian normalizable hypergroup ${(H,\mathcal{B},\mu_{1})}$ is Burnside if and only if the above identity becomes
(4.4) $\prod_{j=1}^{m}\mu_{j}=\sum_{x_{i}\in{G{(H,\mathcal{B})}}}{\mathrm{sgn}}(x_{i}){\widehat{E}}_{i}$
@@
It follows from Proposition 4.1. Recall from Definition 2.34 that ${\mathrm{sgn}}(x_{i})$ is the determinant of $L_{\frac{x_{i}}{d_{i}}}$, when $x_{i}\in{G{(H,\mathcal{B})}}$. ∎
@@
(4.5) $(\prod_{j=1}^{m}\mu_{j})^{2}=\sum_{x_{i}\in{G{(H,\mathcal{B})}}}{\widehat{E}}_{i}.$
@@
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup such that ${G{(H,\mathcal{B})}}$ has odd order. Then ${(H,\mathcal{B})}$ is Burnside if and only if
(4.6) $\prod_{j=1}^{m}\mu_{j}=\sum_{x_{i}\in G(H,B)}{\widehat{E}}_{i}$
@@
Note that in this case ${\mathrm{sgn}}(x_{i})=1$ for all $x_{i}\in{G{(H,\mathcal{B})}}$, since each ${\frac{x_{i}}{d_{i}}}\in{G{(H,\mathcal{B})}}$ has odd order. ∎
@@
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup. We say that the hypergroup ${(H,\mathcal{B})}$ is dual-Burnside if ${\widehat{(H,\mathcal{B})}}$ is Burnside. In other words, for all $\mu_{j}\in{\widehat{\mathcal{B}}}$, the following are equivalent:
@@
The dual of Corollary 4.4 is the following:
@@
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup. Then ${(H,\mathcal{B},\mu_{1})}$ is dual-Burnside if and only if:
(4.7) $\prod_{j=1}^{m}{\frac{x_{i}}{d_{i}}}=\sum_{\mu_{j}\in G{\widehat{(H,\mathcal{B})}}}{\mathrm{sgn}}(\mu_{j})F_{j},$
where ${\mathrm{sgn}}(\mu_{j})$ is the determinant of the left multiplication operator by $\mu_{j}$ on $H^{*}$.
@@
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup. Then ${(H,\mathcal{B},\mu_{1})}$ is dual-Burnside if and only if :
(4.8) $(\prod_{j=1}^{m}{\frac{x_{i}}{d_{i}}})^{2}=\sum_{\mu_{j}\in G{\widehat{(H,\mathcal{B})}}}F_{j}.$
where ${\mathrm{sgn}}(\mu_{j})$ is the determinant of the left multiplication operator by $\mu_{j}$ on $H^{*}$.
@@
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup. If ${{G{\widehat{(H,\mathcal{B})}}}}$ has odd order then ${(H,\mathcal{B})}$ is dual-Burnside if and only if
(4.9) $\prod_{j=1}^{m}{\frac{x_{i}}{d_{i}}}=\sum_{\mu_{j}\in{{G{\widehat{(H,\mathcal{B})}}}}}F_{j}.$
4.1. Inequalities for $P$ and ${\widehat{P}}$

### v11

@@ 3. Galois symmetries for hypergroups
Let $(H,\mathcal{B})$ be an abelian rational hypergroup (i.e. $N^{k}_{ij}\in\mathbb{Q}$). For any standard element $x_{i}\in\mathcal{B}$, we can write $x_{i}=\sum_{j\in\mathcal{I}}{\alpha_{{}_{ij}}}F_{j}$, with ${\alpha_{{}_{ij}}}=\mu_{j}(x_{i})\in\mathbb{C}$. Let $\mathbb{K}=\mathbb{Q}({\alpha_{{}_{ij}}})$ be the field obtained by adjoining all ${\alpha_{{}_{ij}}}$ to $\mathbb{Q}$.
@@ 3.1. Permutation of characters $\mu_{j}$
Let ${(H,\mathcal{B})}$ be any abelian rational hypergroup. For any character $\mu_{j}:H\rightarrow\mathbb{C}$ and $\sigma\in\mathrm{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$, define $\sigma.\mu_{j}\in H^{*}$ as the linear function on $H$ such that $[\sigma.\mu_{j}](x_{i})=\sigma(\mu_{j}(x_{i}))=\sigma({\alpha_{{}_{ij}}})$, for all $x_{i}\in{\mathcal{B}}$.
@@
For any abelian rational hypergroup ${(H,\mathcal{B})}$, the function $\sigma.\mu_{j}$ is an algebra map. Thus $\mathrm{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$ acts on the set $^{\widehat{\mathcal{B}}}$.
@@
Suppose that $x_{i_{1}}x_{i_{2}}=\sum_{k\in\mathcal{I}}N^{k}_{{i_{1}},{i_{2}}}x_{k}$. Since $N^{k}_{{i_{1}},{i_{2}}}\in\mathbb{Q}$,
$[\sigma.\mu_{j}]({x_{i_{1}}}{x_{i_{2}}})=[\sigma.\mu_{j}](\sum_{k\in\mathcal{I}}N^{k}_{{i_{1}},{i_{2}}}x_{k})=\sum_{k\in\mathcal{I}}N^{k}_{{i_{1}},{i_{2}}}[\sigma.\mu_{j}](x_{k})=\sum_{k\in\mathcal{I}}N^{k}_{{i_{1}},{i_{2}}}\sigma(\mu_{j}(x_{k})).$
On the other hand,
@@
$=$ $\sigma(\mu_{j}({x_{i_{1}}}{x_{i_{2}}}))=\sigma(\mu_{j}(\sum_{k\in\mathcal{I}}N^{k}_{{i_{1}},{i_{2}}}x_{k}))$
$=$ $\sum_{k\in\mathcal{I}}N^{k}_{{i_{1}},{i_{2}}}\sigma(\mu_{j}(x_{k})).$
Thus $[\sigma.\mu_{j}]({x_{i_{1}}}{x_{i_{2}}})=[\sigma.\mu_{j}]({x_{i_{1}}})[\sigma.\mu_{j}]({x_{i_{2}}})$, so that $\sigma.\mu_{j}$ is an algebra map. Now,
$[\sigma.(\sigma^{\prime}.\mu_{j})](x_{i})=\sigma([\sigma^{\prime}.\mu_{j}](x_{i}))=\sigma(\sigma^{\prime}(\mu_{j}(x_{i})))=(\sigma\sigma^{\prime})(\mu_{j}(x_{i}))=[\sigma\sigma^{\prime}.\mu_{j}](x_{i})$
It follows that $\mathrm{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$ acts on the set $^{\widehat{\mathcal{B}}}$. ∎
Thus, there is a permutation $\tau_{\sigma}$ of ${\mathcal{I}}$ such that $\sigma.\mu_{j}=\mu_{\tau_{\sigma}(j)}$, and therefore
(3.1) $\sigma({\alpha_{{}_{ij}}})=\alpha_{{}_{i\tau_{\sigma}(j)}}\;\;\text{for all}\;i,j\in{\mathcal{I}}.$
@@
Equation (3.1) shows that $\sigma(\mathbb{K})\subseteq\mathbb{K}$, for all $\sigma\in\mathrm{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$. Therefore, this is a normal extension and consequently a Galois extension (as ${\mathrm{c}har}(\mathbb{Q})=0$). ∎
@@
The map $\rho:\mathrm{Gal}(\mathbb{K}/\mathbb{Q})\rightarrow\mathbb{S}_{\mathcal{I}},\sigma\mapsto\tau_{\sigma}$ is injective. Indeed, if $\tau_{\sigma}=\tau_{\sigma^{\prime}}$, then $\sigma({\alpha_{{}_{ij}}})=\sigma^{\prime}({\alpha_{{}_{ij}}})$, i.e. $\sigma^{-1}\circ\sigma^{\prime}({\alpha_{{}_{ij}}})={\alpha_{{}_{ij}}}$, for all $i,j$. Thus $\sigma=\sigma^{\prime}$.
@@
Let $(H,\mathcal{B})$ be an abelian rational hypergroup. For any $\sigma\in\mathrm{Gal}(\mathbb{K}/\mathbb{Q})$, let $\tau=\tau_{\sigma}$, with the above notations,
(3.2) $\sigma(n_{j})={n_{\tau(j)}}.$
@@
Applying $\sigma\in\mathrm{Gal}(\mathbb{K}/\mathbb{Q})$ to the orthogonality relation (2.5), we obtain:
$\sum_{i\in\mathcal{I}}h_{i}\mu_{\tau(j)}(x_{i})\mu_{\tau(k)}(x_{i^{*}})=\delta_{j,k}\sigma(n_{j}),$
@@
$\sum_{i\in\mathcal{I}}h_{i}\mu_{\tau(j)}(x_{i})\mu_{\tau(k)}(x_{i^{*}})=\delta_{\tau(j),\tau(k)}{n_{\tau(j)}}.$
Therefore, with $j=k$, we get that $\sigma(n_{j})={n_{\tau(j)}}$. ∎
@@
A hypergroup ${(H,\mathcal{B})}$ is $h$-integral if $h_{i}:=1/N^{1}_{i,i^{*}}\in\mathbb{Z}$, for all $i\in\mathcal{I}$.
@@
Let ${(H,\mathcal{B},\mu_{1})}$ be the Grothendieck ring of any weakly-integral fusion category $\mathcal{C}$. By [Ost15, Theorem 2.13] and Equation (2.11), $\widehat{h}_{i}=\frac{n{(H,\mathcal{B},\mu_{1})}}{n_{i}}$ is an integer, for all $i\in\mathcal{I}$. Thus, the dual $^{\widehat{{K(\mathcal{C})}}}$ is $h$-integral. See also Remark 9.1.
@@
Suppose that ${(H,\mathcal{B},\mu_{1})}$ is an abelian normalizable hypergroup with a $h$-integral dual. Then it is weakly-integral. Moreover, if it is rational, then $\widehat{h}_{k}=\widehat{h}_{\tau(k)}$, for all $k\in\mathcal{I}$.
@@
By Lemma 2.8 and Definition 2.5,
$n{(H,\mathcal{B},\mu_{1})}=n{\widehat{{(H,\mathcal{B},\mu_{1})}}}=\sum_{i\in\mathcal{I}}{\widehat{h}}_{i}\in\mathbb{Z}.$
If ${(H,\mathcal{B},\mu_{1})}$ is rational, then Equation (3.2) can be written as $\sigma\big{(}\frac{n{(H,\mathcal{B},\mu_{1})}}{\widehat{h}_{i}}\big{)}=\frac{n{(H,\mathcal{B},\mu_{1})}}{\widehat{h}_{\tau(i)}}.$ In particular, $\widehat{h}_{k}=\widehat{h}_{\tau(i)}$, since $^{\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is $h$-integral. ∎
@@ 4. Burnside’s vanishing property for hypergroups
Let $\mathcal{B}_{0}$ be the set of all vanishing elements of $\mathcal{B}$, and $\mathcal{B}_{1}$ be the set of all non-vanishing elements of $\mathcal{B}$. Thus $\mathcal{B}=\mathcal{B}_{1}\sqcup\mathcal{B}_{0}$. By the proof of Lemma 2.27, if ${(H,\mathcal{B})}$ is a RN-hypergroup then ${G{(H,\mathcal{B})}}\subseteq\mathcal{B}_{1}$. In the abelian case, since $x_{i}=\sum_{j\in\mathcal{I}}\mu_{j}(x_{i})F_{j}$, it follows that $x_{i}$ is non-vanishing if and only if it is invertible in $H$ (i.e. $\det(L_{x_{i}})$ nonzero). Thus, in this case, $\mathcal{B}_{1}$ coincides with the set of all standard elements that are invertible in $H$.
In any abelian normalizable hypergroup ${(H,\mathcal{B},\mu_{1})}$:
(4.1) ${\widehat{P}}:=\prod_{j\in\mathcal{I}}\mu_{j}=\sum_{i\in\mathcal{I}}{\widehat{P}}({\frac{x_{i}}{d_{i}}})\widehat{F}_{i}.$
@@
Let ${(H,\mathcal{B},\mu_{1})}$ be any abelian normalizable hypergroup. Then:
(4.2) $\prod_{j\in\mathcal{I}}\mu_{j}=\sum_{x_{i}\in\mathcal{B}_{1}}\det(L_{{\frac{x_{i}}{d_{i}}}})\widehat{F}_{i}$
@@
Recall that $\widehat{F}_{i}(\frac{x_{j}}{d_{j}})=\delta_{i,j}$, and $\frac{x_{j}}{d_{j}}=\sum_{i\in\mathcal{I}}\mu_{i}(\frac{x_{j}}{d_{j}})F_{i}$. Thus, for all $j\in\mathcal{I}$,
$\left[\sum_{i\in\mathcal{I}}\det(L_{{\frac{x_{i}}{d_{i}}}})\widehat{F}_{i}\right](\frac{x_{j}}{d_{j}})=\det(L_{\frac{x_{j}}{d_{j}}})=\prod_{i\in\mathcal{I}}\mu_{i}(\frac{x_{j}}{d_{j}})=\left[\prod_{i\in\mathcal{I}}\mu_{i}\right](\frac{x_{j}}{d_{j}})$
It follows that
$\prod_{i\in\mathcal{I}}\mu_{i}=\sum_{i\in\mathcal{I}}\det(L_{{\frac{x_{i}}{d_{i}}}})\widehat{F}_{i}=\sum_{x_{i}\in\mathcal{B}_{1}}\det(L_{{\frac{x_{i}}{d_{i}}}})\widehat{F}_{i}.$
Indeed, $x_{i}\in{\mathcal{B}_{1}}$ if and only if $\det(L_{{\frac{x_{i}}{d_{i}}}})$ is nonzero, justifying the last equality.∎
Dually, applying the above result to $^{\widehat{{(H,\mathcal{B},\mu_{1})}}}$, we obtain a decomposition of the basis ^=^${\widehat{\mathcal{B}}}={\widehat{\mathcal{B}}}_{0}\sqcup{\widehat{\mathcal{B}}}_{1}$. Proposition 4.1 implies the following:
@@
Let ${(H,\mathcal{B},\mu_{1})}$ be any abelian normalizable hypergroup. Then:
(4.3) $\prod_{i\in\mathcal{I}}{\frac{x_{i}}{d_{i}}}=\sum_{\mu_{j}\in{\widehat{\mathcal{B}}}_{1}}\det(L_{\mu_{j}})F_{j},$
@@
An abelian hypergroup ${(H,\mathcal{B})}$ is Burnside if and only if we have ${G{(H,\mathcal{B})}}=\mathcal{B}_{1}$.
@@
An abelian normalizable hypergroup ${(H,\mathcal{B},\mu_{1})}$ is Burnside if and only if the following equality holds:
(4.4) $\prod_{j\in\mathcal{I}}\mu_{j}=\sum_{x_{i}\in{G{(H,\mathcal{B})}}}{\mathrm{sgn}}(x_{i})\widehat{F}_{i},$
where the notation ${\mathrm{sgn}}(x_{i})$ was introduced in Definition 2.41.
@@
Immediate from Propositions 4.1 and 4.3. ∎
@@
(4.5) $(\prod_{j\in\mathcal{I}}\mu_{j})^{2}=\sum_{x_{i}\in{G{(H,\mathcal{B})}}}\widehat{F}_{i}.$
@@
An abelian normalizable hypergroup ${(H,\mathcal{B},\mu_{1})}$, such that ${G{(H,\mathcal{B})}}$ has odd order, is Burnside if and only if
(4.6) $\prod_{j\in\mathcal{I}}\mu_{j}=\sum_{x_{i}\in G(H,B)}\widehat{F}_{i}.$
@@
Every ${\frac{x_{i}}{d_{i}}}\in{G{(H,\mathcal{B})}}$ has odd order, so ${\mathrm{sgn}}(x_{i})=1$. ∎
@@
An abelian normalizable hypergroup ${(H,\mathcal{B},\mu_{1})}$ is called dual-Burnside if $^{\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is Burnside; in other words, if for all $\mu_{j}\in{\widehat{\mathcal{B}}}$, the following are equivalent:
@@
This can be shortened to $G\widehat{{(H,\mathcal{B})}}={\widehat{\mathcal{B}}}_{1}$.
Here are the dual of Corollaries 4.4, 4.5 and 4.6:
@@
An abelian normalizable hypergroup ${(H,\mathcal{B},\mu_{1})}$ is dual-Burnside if and only if:
(4.7) $\prod_{i\in\mathcal{I}}{\frac{x_{i}}{d_{i}}}=\sum_{\mu_{j}\in G{\widehat{(H,\mathcal{B})}}}{\mathrm{sgn}}(\mu_{j})F_{j}.$
@@
An abelian normalizable hypergroup ${(H,\mathcal{B},\mu_{1})}$ is dual-Burnside if and only if :
(4.8) $(\prod_{i\in\mathcal{I}}{\frac{x_{i}}{d_{i}}})^{2}=\sum_{\mu_{j}\in G{\widehat{(H,\mathcal{B})}}}F_{j}.$
@@
An abelian normalizable hypergroup ${(H,\mathcal{B},\mu_{1})}$, with $^{{G{\widehat{(H,\mathcal{B})}}}}$ of odd order, is dual-Burnside if and only if
(4.9) $\prod_{i\in\mathcal{I}}{\frac{x_{i}}{d_{i}}}=\sum_{\mu_{j}\in{{G{\widehat{(H,\mathcal{B})}}}}}F_{j}.$
4.1. Inequalities for $P$ and $^{\widehat{P}}$
Recall that $P:=\prod_{i\in\mathcal{I}}\frac{x_{i}}{d_{i}}$ and ${\widehat{P}}:=\prod_{j\in\mathcal{I}}\mu_{j}$.